The sum of two numbers is 135. The difference between them is 9. Find the numbers.
Answer by nerdybill(7384)

The sum of two numbers is 135. The difference between them is 9. Find the numbers.
.
Let x = first number
and y = second number
.
from: "The sum of two numbers is 135" we get
x + y = 135 (equation 1)
.
from: "The difference between them is 9" we get
x - y = 9 (equation 2)
.
Solve equation 2 for y:
x - y = 9
x - 9 = y
.
Substitute the above into equation 1 and solve for x:
x + y = 135
x + x - 9 = 135
2x - 9 = 135
2x = 144
x = 72
.
Substitute above into equation 2 and solve for y:
x - y = 9
72 - y = 9
-y = -63
y = 63
.
Answer: 63 and 72